OpenStudy (anonymous):

Can someone please give a basic definition/example of a parabola, circle, ellipse and hyperbola equation.

OpenStudy (anonymous):

A parabola is basically a curve shape. The dot in the middle represents the focus and the shape around it like a "U" shape is the parabola itself. 

OpenStudy (anonymous):

ok, and what would an equation for this look like?

OpenStudy (anonymous):

The circle is a basic shape. If you know how to find the area of a circle, then you should know what it is. If not this is an example: 

OpenStudy (anonymous):

I quite forgot what's the formula for the Parabola.

OpenStudy (anonymous):

However, for the parabola, its finding the coordinates.

OpenStudy (anonymous):

@XxINFINITYxX you're with me so far?

OpenStudy (anonymous):

yep, sorry I'm helping someone else out at the same time, and i understand what the actual graph will look like, i was just confused about what the equation differentiation between all of them is

OpenStudy (anonymous):

Formula: Parabola - not sure what's the formula or if it has one. Circle: Circumference - C=2(pi)*r Radius - r=C/2(pi) Area - A=(pi)r^2 Ellipse - TF1 + TF2 = K Hyperbola - x^2/a^2-y^2/b^2=1

OpenStudy (anonymous):

@XxINFINITYxX

OpenStudy (anonymous):

ok that helped, so what if I had an equation like: (1/80)*y^2=x what type of graph would that be? a circle?

OpenStudy (anonymous):

I'm sorry, but it has been a while that I have not solved a parabola. I'm kinda rusty at the moment.

OpenStudy (anonymous):

its okay! thanks for your time and effort

OpenStudy (anonymous):

Do you need me to show you the rest or you're good?

OpenStudy (anonymous):

I should be ok, I think I can figure it out on my own, but thanks anyway

OpenStudy (anonymous):

Although, I did send you the formula for it - so that might help you.